What volume of 3 molar HN0_3  is needed to oxidise of 8g of Fe^{2+} to Fe^{3+}? HNO_3 get converted to NO :

a) 

8 ml

b) 

16 ml

c) 

32 ml

d) 

64 ml
 

Grade:12

2 Answers

Arun
4 years ago

It is defined as the addition of oxygen / electronegative element to a substance or removal of hydrogen / electropositive element from a substance.

- wherein

e.g. 2H_{2}S(g) + O_{2}(g)\rightarrow 2S(s)+2H_{2}O(l)

 

 \underset{77g}{HNO_{3}}+\underset{168g}{3Fe^{2+}}\rightarrow 3Fe^{3+}+side\, \,products

Now 168g of Fe^{2+}\rightarrow 77g\, \, HNO_{3}

8gFe^{2+}\rightarrow \frac{77\times 8}{168}=3.67g

Molarity (given)=3M

3=\frac{w\times 1000}{mV}

V=\frac{3.67\times 1000}{mV}

V=\frac{3.67\times 1000}{3\times 77}= 15.88ml\sim 16ml

Khimraj
4 years ago
HNO3 + 3 Fe2+ --------> 3 Fe3+ +side products
77 g 168 g 
168 g of Fe2+ forms HNO3 = 77g
8 g of Fe2+ forms HNO3 = 77x8/168
= 3.67 g
molarity = 3M
molarity = w x 1000/Mx V
3 = 3.67 x 1000/77x V
V= 15.88 ml